# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v2

@@ V Limitation and Future Work
Although our method achieves outstanding results, it also has some limitations. First, since MAL-SPC relies on the geometric similarity of objects within the same category, it requires category information of the partial observations. Besides, our model performs poorly on unseen categories and rare objects, despite outperforming other self-supervised approaches. In the future, we will attempt to eliminate the dependency on category information using methods such as clustering and contrastive learning. Additionally, we plan to leverage large quantities of 2D images and current Large Vision Models to provide 2D priors for partial point clouds, so as to improve the generalization capability of self-supervised methods.
@@ VI Conclusion
In the scenario of self-supervised point cloud completion, it is critical to fully utilize the prior geometries from partial observations due to the lack of complete shapes as supervision. To this end, we have presented a novel approach that effectively exploits the objects’ intrinsic region-level and category-specific geometric similarity to infer complete shapes without the need for multi-view partial scans and masked operation. Our method demonstrates strong capabilities in both inferring complete shapes and preserving incomplete input, resulting in reconstructed shapes that are both reasonable and accurate, with minimal noise. Quantitatively and visually, our approach surpasses current state-of-the-art self-supervised methods, as well as a portion of unpaired methods. We hope this work inspires future research in 3D object and scene completion, reconstruction, and understanding tasks.

### v3

@@ V Limitation and Future Work
Although our method achieves outstanding results, it also has some limitations. First, since MAL-UPC relies on the geometric similarity of objects within the same category, it requires category labels to allocate the raw depth maps of the partial point clouds to their respective category-specific image banks. Besides, our model performs poorly on unseen categories and rare objects, despite outperforming self-supervised approaches. In the future, we will attempt to eliminate the dependency on category labels using methods such as clustering and contrastive learning. Additionally, we plan to leverage large quantities of 2D images and current Large Vision Models to provide 2D priors for partial point clouds, so as to improve the generalization capability of unsupervised methods.
@@ VI Conclusion
In the scenario of unsupervised point cloud completion, it is critical to fully utilize the prior geometries from partial observations due to the lack of complete shapes as supervision. To this end, we have presented a novel approach that effectively exploits the objects’ intrinsic region-level and category-specific geometric similarity to infer complete shapes without the need for multi-view partial scans and masked operation. Our method demonstrates strong capabilities in both inferring complete shapes and preserving incomplete input, resulting in reconstructed shapes that are both reasonable and accurate, with minimal noise. Quantitatively and visually, our approach surpasses current state-of-the-art self-supervised methods, as well as a portion of unpaired methods. We hope this work inspires future research in 3D object and scene completion, reconstruction, and understanding tasks.